22=6(2x-3)-3(5x-8)
how do I find what x is?

Answer by ikdeep(226)

22=6(2x-3)-3(5x-8)
1st starat with the barckets....
22= (12x - 18)- (15x - 24)
22 = 12x - 18 -15x + 24 (you must have the idea that if before brackets there is a "-" sign then when we open the brackets all the signs are changed)
Now, put together the like variables,,,
22 = 12x -15x - 18 + 24
22 = -3x + 6
22 -6 = -3x
16 = -3x
16 / -3 = x
Or -16/3 = x